A metal of length 15.0lm is heated until temperature rise by 60c° If it's New length is 15.05m.calculate The linear expensivity




1
Expert's answer
2022-03-01T17:33:36-0500

By definition, the linear expensivity is:

"\alpha = \dfrac{1}{L}\dfrac{\Delta L}{\Delta T} = \dfrac{1}{15.01m}\cdot \dfrac{(15.05m-15.01m)}{60\degree C} \approx 4.4\times 10^{-5}\degree C^{-1}"

Answer. "4.4\times 10^{-5}\degree C^{-1}".